OpenStudy (anonymous):

help with...4x^2+8x-21=0

myininaya (myininaya):

do you know the quadratic formula?

OpenStudy (anonymous):

no..i don think so?

myininaya (myininaya):

ok we can try to factor it then

OpenStudy (anonymous):

yeah thats what im doing..i have to factor it

myininaya (myininaya):

a=4, b=8, c=-21 Find factors of a*c that add up to be b

myininaya (myininaya):

what is a*c?

OpenStudy (anonymous):

a*c=-84

myininaya (myininaya):

right! a*c=-4(21) play with this until you get the sum of the factors to be b=8 -4+21 is not 8 a*c=12(-7) 12-7 is not 8 a*c=6(-14) is 6-14 =8?

OpenStudy (anonymous):

-8 yeah

myininaya (myininaya):

but we want 8 how can we get 8

myininaya (myininaya):

what if you said a*c=-6(14) what is the sum of those two factors?

OpenStudy (anonymous):

...8!

myininaya (myininaya):

right!

myininaya (myininaya):

now replace the bx term with what you just found but you need an x behind each term so it equals bx

myininaya (myininaya):

4x^2+14x-6x-21=0

OpenStudy (anonymous):

ohhh okay then you have to group it right?

myininaya (myininaya):

right!

OpenStudy (anonymous):

so then it would be(4x^2+14x)(-6x-21)=0 which then what?

myininaya (myininaya):

no!

OpenStudy (anonymous):

what??

myininaya (myininaya):

what do 4x^2 and 14x have in common?____________ what do -6x and -21 have in common?___________

OpenStudy (anonymous):

4x ? idk

myininaya (myininaya):

2x -3 so the first 2 terms we factor out 2x(2x+7) so the second 2 terms we factor out -3(2x+7) so 4x^2+14x-6x-21=2x(2x+7)-3(2x+7) can we do anything else?

myininaya (myininaya):

you still there?

myininaya (myininaya):

we are almost there

OpenStudy (anonymous):

yeha ohh aha i would have never thought of that..

OpenStudy (anonymous):

then you do 2x-3=0 and 2x+7=0..

myininaya (myininaya):

there is 2x+7 in both of those two terms above so we can factor a 2x+7 out of both giving us (2x+7)(2x-3) this expression was orignally set =0 so we have (2x+7)(2x-3)=0 so either 2x+7=0 or 2x-3=0 when i say both its the exclusive kind

myininaya (myininaya):

right

myininaya (myininaya):

you got that part? was the factoring the problem you were having?

OpenStudy (anonymous):

then the answer is x=2/3 and x=2/7!!

OpenStudy (anonymous):

omg thanks soo much for your help!! haha sorry that took me forever to figure out:/ thanks though!

myininaya (myininaya):

x=2/3 is right but x=-7/2 for the second one